# Teleportation=Translation: Continuous recovery of black hole information

### v3

@@ 3.2 Path A: Pedagogical (CP-Map) Interpolation
To see this failure explicitly, we refer the reader to the $M_{2}(\mathbb{C})$ toy model in Appendix B, where the naive convex-combination path (Path A) is constructed in a finite-dimensional setting. There, a direct computation verifies the breakdown of idempotency, even in this simple Type I context. This counterexample illustrates that the defect is not specific to the infinite-dimensional setting but is intrinsic to the interpolation strategy itself. It underscores the need for a more principled and mathematically rigorous approach—namely, the Haagerup–Kosaki construction (Path B) introduced in the next section.
@@ 3.3 Path B: The Canonical Haagerup–Kosaki Interpolation
Ideally, to resolve the idempotency failure identified in Path A, one would construct a continuous family of genuine conditional expectations. However, the existence of such trace-preserving expectations is obstructed by the intrinsic nature of the physical algebra. In physical settings such as algebraic quantum field theory, the relevant operator algebras–typically denoted $\mathcal{M}$ and $\mathcal{N}$–are of Type III. A hallmark of such algebras is the absence of a normal, faithful, tracial state. Physically, this reflects the fact that the vacuum state exhibits unbounded entanglement across spatial regions, rendering it impossible to assign finite information measures in the usual way. Mathematically, this precludes the existence of a trace-preserving conditional expectation $E:\mathcal{M}\rightarrow\mathcal{N}$. Instead, one must rely on a faithful normal operator-valued weight (OVW) $\mathcal{E}:\mathcal{M}^{+}\rightarrow\widehat{\mathcal{N}}^{+}$, which lacks the boundedness and trace properties required for our purposes. Here, $\mathcal{M}^{+}$ and $\mathcal{N}^{+}$ are the positive cones of the respective algebras, while $\widehat{\mathcal{N}}^{+}$ denotes the extended positive cone of $\mathcal{N}$, a domain that mathematically accommodates the infinite values (UV divergences) inherent in the Type III setting.
To resolve this structural deficit, we employ the continuous crossed-product construction, a cornerstone of the Haagerup–Kosaki framework. This procedure embeds the original Type III algebra $\mathcal{M}$ into a larger semifinite envelope $\tilde{\mathcal{M}}$ (typically of Type $\mathrm{II}_{\infty}$) via Takesaki duality. Crucially, this dual algebra admits a canonical semifinite trace $\tau$, which provides the necessary tracial background absent in the physical algebra. The connection between the envelope and the base algebra is formally established by the canonical structural map $\underline{\mathcal{E}}:\tilde{\mathcal{M}}^{+}\to\widehat$$\mathcal{M}$$^{+}$. As the fundamental operator-valued weight of the crossed-product, $\underline{\mathcal{E}}$ serves as a structural bridge that relates the canonical trace $\tau$ to the physical weights on $\mathcal{M}$.
Within this lifted setting, the original inclusion $\mathcal{N}\subset\mathcal{M}$ and its associated OVW $\mathcal{E}$ are regularized into a genuine, trace-preserving conditional expectation $\tilde{E}:\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{N}}$. While the structural weight $\underline{\mathcal{E}}$ ensures the mathematical consistency of the lifting itself, the conditional expectation $\tilde{E}$ provides the idempotent projection required to define a valid interpolation path. Following the insight of Kosaki’s theorem, the existence of such a trace-preserving projection allows the static inclusion to be dynamically resolved: it enables the construction of a continuous family of maps connecting the identity on $\tilde{\mathcal{M}}$ to the projection $\tilde{E}$ on $\tilde{\mathcal{N}}$. This mathematical mechanism transforms the algebraic inclusion into a continuous flow, providing the foundation for our information recovery protocol.
@@
The Haagerup–Kosaki theory TakesakiII ; Kosaki84 implements this by constructing the crossed-product algebra using a faithful normal semifinite weight $\omega$ on $\mathcal{M}$ that is chosen to be compatible with the inclusion (i.e., $\omega=\psi\circ\mathcal{E}$ for some weight $\psi$ on $\mathcal{N}$). Note that the existence of such a weight is guaranteed by Haagerup’s theory, and this choice does not result in any loss of generality regarding the structure of the crossed product. Using its associated modular automorphism group $\sigma^{\omega}$, which leaves $\mathcal{N}$ invariant, we construct the crossed-product algebras:
In this larger, tracial semifinite-envelope algebra ($\tilde{\mathcal{M}}$ equipped with a faithful semifinite trace $\tau$), the original OVW $\mathcal{E}$ is realized as a genuine, faithful normal conditional expectation $\tilde{E}:\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{N}}$. This transformation from a non-tracial to a tracial setting facilitates the subsequent construction of the physical idempotent path. The formal statement and the necessary conditions for the existence of $\tilde{E}$ are detailed in Appendix C.
Throughout this paper, we employ the tilde notation (e.g., $\tilde{\mathcal{M}}$, $\tilde{E}$) to distinguish objects in this lifted semifinite-envelope algebra from their original Type III counterparts (e.g., $\mathcal{M}$, $\mathcal{E}$). This algebra $\tilde{\mathcal{M}}$ serves as the workspace where our well-defined, continuous path is constructed.
However, the existence of this lifted structure does not, by itself, guarantee a valid interpolation path. One might naively attempt to construct intermediate algebras via spectral projections (or spectral cuts) of the modular operator. While intuitively appealing, such ad-hoc truncations generally fail to preserve the subalgebra structure and, more critically, exhibit pathological boundary behavior. Specifically, any finite spectral cut excludes the tails of the modular spectrum, leaving the algebra effectively open and unable to recover the identity in the $s\to 0$ limit. Consequently, no matter how the limit is taken, such paths cannot close the gap to the full algebra $\tilde{\mathcal{M}}$, resulting in a fundamental discontinuity at $s=0$.
To resolve these boundary issues—specifically to enforce the condition $\tilde{\mathcal{N}}(0)=\tilde{\mathcal{M}}$—we adopt the canonical $L^{p}$-interpolation path derived from the theory of non-commutative $L^{p}$ spaces Kosaki84 . Unlike arbitrary spectral cuts, this framework provides a rigorous method to interpolate between the algebra $\tilde{\mathcal{M}}$ (identified with $L^{\infty}$) and its predual space of normal functionals (identified with $L^{1}$).
Intuitively, the interpolation parameter $p\in[1,\infty]$ connects these spaces, where the intermediate elements behave like fractional powers of states and operators. In our specific context, we reparametrize this interpolation using $s=1/p\in[0,1]$. Crucially, by interpolating between the reference weight $\tilde{\phi}_{0}$ and its restriction $\tilde{\phi}_{1}=\tilde{\phi}_{0}\circ\tilde{E}$, we model the continuous loss of information. This transition follows the modular flow, ensuring that the non-commutative structure is preserved via analytic continuation, providing the unique rigidity required for our canonical path. Based on this structure, we define the interpolation path as follows.
@@
Let $\tilde{\phi}_{0}$ be the faithful normal semifinite dual weight on the crossed product algebra $\tilde{\mathcal{M}}$, constructed from the physical weight $\omega$ on $\mathcal{M}$. We designate $\tilde{\phi}_{0}$ as the reference weight. [^1] Note that $\tilde{\phi}_{0}$ is distinct from the canonical trace $\tau$; while $\tau$ provides the structural background for the lifted setting, $\tilde{\phi}_{0}$ represents the physical reference weight.
We define the endpoint weight $\tilde{\phi}_{1}:=\tilde{\phi}_{0}\circ\tilde{E}$, which represents the weight restricted to the target subalgebra $\tilde{\mathcal{N}}$. This relation implies that $\tilde{\phi}_{1}$ is not an independent state, but rather a coarse-grained version of $\tilde{\phi}_{0}$ filtered through $\tilde{E}$. It also formally lifts the compatibility condition $\omega=\psi\circ\mathcal{E}$ to the level of dual weights on the crossed product. Crucially, whereas $\omega$ ensures the inclusion property during the lifting process, $\tilde{\phi}_{1}$ serves as the target weight on $\tilde{\mathcal{N}}$ that governs the interpolation trajectory.
For $s\in[0,1]$, the canonical $L^{p}$-interpolation constructs a one-parameter family of conditional expectations $\tilde{E}_{s}:\tilde{\mathcal{M}}\to\tilde{\mathcal{M}}$. The corresponding algebras $\tilde{\mathcal{N}}(s)$ are defined as the range of these expectations:
@@
The intermediate expectations $\tilde{E}_{s}$ are uniquely determined by the analytic continuation of the Connes–Takesaki Radon–Nikodym cocycle $D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}$. Specifically, if $h=d\tilde{\phi}_{1}/d\tilde{\phi}_{0}$ denotes the non-commutative Radon–Nikodym derivative, the interpolation is driven by the relation $D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}=h^{it}$. This relation underpins the analyticity of the Haagerup $L^{p}$-interpolation path at the boundaries (particularly at $s=0$) and ensures that the interpolated weights $\tilde{\phi}_{s}$ are uniquely determined by the resulting non-commutative $L^{p}$ space structure.
The construction in Definition 3.5 has a clear physical interpretation in terms of non-commutative geometry. The Radon–Nikodym derivative relation implies that the interpolated weight $\tilde{\phi}_{s}$ satisfies the exact power-law relation regarding its density:
Heuristically, this signifies that $\tilde{\phi}_{s}$ acts as a non-commutative geometric mean between the reference weight $\tilde{\phi}_{0}$ and the coarse-grained weight $\tilde{\phi}_{1}$. Although formal products of weights are not defined in the operator algebra, we may conceptually visualize this interpolation as:
@@
This geometric averaging property ensures that the path $\tilde{\mathcal{N}}(s)$ follows the geodesic of the underlying modular structure, minimizing the information distance between the full algebra and the subalgebra.
This construction naturally yields the corresponding conditional expectations $\tilde{E}_{s}:\tilde{\mathcal{M}}\to\tilde{\mathcal{N}}(s)$. We now state the fundamental properties of this path, analogous to the properties we sought in the naive constructions.
Theorem 3.6 (Properties of the Canonical Path).
@@
    Nesting (Monotonicity): For $s^{\prime}\geq s$, we have the inclusion $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$. This reflects the monotonic coarse-graining nature of the flow, analogous to a renormalization group flow.
@@
These properties follow directly from the analytic structure of the Haagerup-Kosaki interpolation. Unlike naive spectral truncations, this path relies on complex interpolation between the weights $\tilde{\phi}_{0}$ and $\tilde{\phi}_{1}$, ensuring structural continuity via the modular automorphism group.
(1) Boundary Conditions: The interpolation relies on the analytic continuation of the Radon-Nikodym derivative $h$ defined above. At $s=0$, the interpolation yields the unperturbed reference weight $\tilde{\phi}_{0}$, implying that the associated projection is the identity map ($\tilde{E}_{0}=\text{id}_{\tilde{\mathcal{M}}}$). Conversely, at $s=1$, the construction converges in norm to the restricted weight $\tilde{\phi}_{1}$, thereby recovering the original conditional expectation $\tilde{E}_{1}=\tilde{E}$ and its range $\tilde{\mathcal{N}}$. This smooth limit behavior eliminates the discontinuity at $s=0$ that plagues spectral cut methods.
(2) Nesting: This property arises from the definition $\tilde{\phi}_{1}=\tilde{\phi}_{0}\circ\tilde{E}$. The operator $h$ essentially acts as a density operator representing the projection $\tilde{E}$. The interpolation corresponds to powers $h^{s}$, which serve as a soft filter whose intensity increases with $s$. Since $0\leq h\leq 1$ (in the appropriate operator sense), increasing $s$ strictly contracts the support of the associated expectations. Thus, for $s^{\prime}\geq s$, the range of $\tilde{E}_{s^{\prime}}$ is contained within the range of $\tilde{E}_{s}$.
(3) Idempotency: This follows directly from the nesting property. Since $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ for $s^{\prime}\geq s$, the operator $\tilde{E}_{s^{\prime}}$ projects onto a subspace contained within the range of $\tilde{E}_{s}$. Thus, applying the coarser filter $\tilde{E}_{s}$ before the finer filter $\tilde{E}_{s^{\prime}}$ leaves the latter invariant: $\tilde{E}_{s^{\prime}}\tilde{E}_{s}=\tilde{E}_{s^{\prime}}$. ∎
The canonical $L^{p}$-interpolation provides more than mathematical smoothness; it imposes a strict ordering on information. The nesting property $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ established in Theorem 3.6 identifies $s$ as a coarse-graining scale: increasing $s$ corresponds to systematically discarding information, strictly analogous to the flow of a renormalization group. Conversely, the unitary operator $\tilde{U}(s)$ generated by this flow acts as an information recovery process, reversing this coarse-graining by translating information from the finer-grained algebra back into the decodable sector. Thus, this construction not only connects the two algebras but also establishes the rigorous framework required for the unitary recovery path.

### v4

@@ 3.2 Path A: Pedagogical Interpolation
To illustrate this failure explicitly, we refer the reader to the $M_{2}(\mathbb{C})$ toy model in Appendix B. There, a direct computation verifies the breakdown of idempotency for the linear path, even in the simplest finite-dimensional Type I context. This counterexample demonstrates that the defect is not an artifact of the infinite-dimensional setting but is intrinsic to the linear interpolation strategy itself. This observation underscores the necessity for the strictly algebraic, modular-covariant approach—namely, the Haagerup–Kosaki construction (Path B)—which we develop in the following section.
@@ 3.3 Path B: The Canonical Haagerup–Kosaki Interpolation
As demonstrated by our analysis of Path A, a physically consistent interpolation must strictly preserve the von Neumann algebra structure throughout the process. This requirement necessitates a continuous family of maps satisfying dynamic idempotency, implying that every intermediate map must be a genuine conditional expectation. However, constructing such a trace-preserving family faces a fundamental obstruction due to the intrinsic nature of the physical algebra.
@@
In the context of algebraic quantum field theory, the local operator algebras $\mathcal{M}$ and $\mathcal{N}$ (where $\mathcal{N}\subset\mathcal{M}$) are of Type III. A defining characteristic of Type III algebras is the absence of a normal, faithful, tracial state. Physically, this reflects the unbounded entanglement of the vacuum, which prevents the assignment of finite local density matrices or information measures. Mathematically, this precludes the existence of a trace-preserving conditional expectation $E:\mathcal{M}\rightarrow\mathcal{N}$. Instead, the relationship between the algebras is governed by a faithful normal operator-valued weight (OVW) $\mathcal{E}:\mathcal{M}^{+}\rightarrow\widehat{\mathcal{N}}^{+}$. This weight is unbounded and lacks the necessary properties to define a normalized projection. Here, $\mathcal{M}^{+}$ and $\mathcal{N}^{+}$ denote the positive cones of the respective algebras, while $\widehat{\mathcal{N}}^{+}$ denotes the extended positive cone of $\mathcal{N}$, a domain that mathematically accommodates the infinite values (UV divergences) inherent in the Type III setting.
To resolve this structural obstruction, we employ the Haagerup-Kosaki framework to lift the entire inclusion to a semifinite setting where a trace exists. We achieve this by constructing the crossed product with respect to a faithful normal semifinite weight $\omega$ on $\mathcal{M}$ chosen to be compatible with the inclusion. Compatibility requires that the modular automorphism group of the weight leaves the subalgebra invariant (i.e., $\sigma_{t}^{\omega}(\mathcal{N})=\mathcal{N}$, implying $\omega=\psi\circ\mathcal{E}$ for some weight $\psi$ on $\mathcal{N}$). Using the associated modular automorphism group $\sigma^{\omega}$, we construct the crossed-product envelopes:
Throughout this paper, we employ the tilde notation (e.g., $\tilde{\mathcal{M}}$, $\tilde{E}$) to distinguish objects in this lifted semifinite-envelope algebra from their original Type III counterparts (e.g., $\mathcal{M}$, $\mathcal{E}$). In this lifted setting, both $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$ become Type II von Neumann algebras equipped with a canonical semifinite trace $\tau$. Within this tracial framework, the original unbounded operator-valued weight $\mathcal{E}$ is regularized into a genuine, trace-preserving conditional expectation $\tilde{E}:\tilde{\mathcal{M}}\rightarrow\tilde{\mathcal{N}}$. This transformation from a non-tracial to a tracial setting facilitates the subsequent construction of the physical idempotent path. The formal statement and necessary conditions for the existence of $\tilde{E}$ are detailed in Appendix C.
@@
However, the existence of this lifted structure does not, by itself, guarantee a valid interpolation path. One might naively attempt to construct intermediate algebras via spectral projections (or spectral cuts) of the modular operator. While intuitively appealing, such ad hoc truncations generally fail to preserve the subalgebra structure and, more critically, exhibit pathological boundary behavior. Specifically, any finite spectral cut excludes the tails of the modular spectrum, leaving the algebra effectively open and unable to recover the identity in the $s\to 0$ limit. Consequently, no matter how the limit is taken, such paths cannot close the gap to the full algebra $\tilde{\mathcal{M}}$, resulting in a fundamental discontinuity at $s=0$.
We construct the continuous interpolation path by bridging these two lifted algebras, $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$, via the analytic structure of non-commutative $L^{p}$ spaces. The key mechanism is to identify the interpolation parameter $s\in[0,1]$ with the $L^{p}$-space parameter $p$ via the relation $s=1/p$. This parameterization physically governs the information resolution of the algebra along the path. Equivalently, by interpolating between the reference weight $\tilde{\phi}_{0}$ and its restriction $\tilde{\phi}_{1}=\tilde{\phi}_{0}\circ\tilde{E}$, we model the continuous loss of information. This transition follows the modular flow, ensuring that the non-commutative structure is preserved via analytic continuation, providing the unique rigidity required for our canonical path.
  • At $s=0$ ($p=\infty$): The system corresponds to the non-commutative $L^{\infty}$-space, which is the algebra itself. Thus, we retain the full information of the envelope algebra $\tilde{\mathcal{M}}$.
  • At $s=1$ ($p=1$): The system corresponds to the non-commutative $L^{1}$-space (the predual), restricted to the subalgebra. Here, the information is fully coarse-grained to the target algebra $\tilde{\mathcal{N}}$.
  • For $0<s<1$: The system possesses an intermediate information resolution, defined by the analytic interpolation between the full algebra $\tilde{\mathcal{M}}$ and the subalgebra $\tilde{\mathcal{N}}$.
Based on this structure, we define the interpolation path as follows.
@@
Let $\tilde{\phi}_{0}$ be the faithful normal semifinite dual weight on the crossed product algebra $\tilde{\mathcal{M}}$, constructed from the physical weight $\omega$ on $\mathcal{M}$. This dual weight $\tilde{\phi}_{0}$ serves as the reference weight for the interpolation.
We define the endpoint weight $\tilde{\phi}_{1}:=\tilde{\phi}_{0}\circ\tilde{E}$, representing the reference weight coarse-grained to the target subalgebra $\tilde{\mathcal{N}}$. Mathematically, this definition lifts the compatibility condition of the original inclusion ($\omega=\psi\circ\mathcal{E}$) to the crossed product level. Crucially, whereas $\tilde{\phi}_{0}$ retains full information, $\tilde{\phi}_{1}$ serves as the target state that governs the trajectory of information loss.
The canonical interpolation is generated by the analytic continuation of the relative modular flow between these two weights. Specifically, the Connes–Takesaki Radon–Nikodym cocycle $D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}$ is generated by the non-commutative Radon–Nikodym derivative $h=d\tilde{\phi}_{1}/d\tilde{\phi}_{0}$. This positive self-adjoint operator induces the cocycle via the relation:
$D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}=h^{it}.$ (13)
We define the interpolated weight $\tilde{\phi}_{s}$ for $s\in[0,1]$ by analytically extending the modular time $t$ to imaginary time (corresponding to the Wick rotation $t\to-is$). This yields the specific density relation:
Since the entire path is generated by the spectral resolution of the single operator $h$, all relative Radon-Nikodym derivatives commute. Consequently, the chain rule for the Connes-Takesaki cocycle simplifies to the algebraic sum of exponents:
$[D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}$ $=[D\tilde{\phi}_{1}:D\tilde{\phi}_{s}]_{t}[D\tilde{\phi}_{s}:D\tilde{\phi}_{0}]_{t}$
$=h^{i(1-s)t}h^{ist}=h^{it}~.$ (15)
Finally, the one-parameter family of algebras $\tilde{\mathcal{N}}(s)$ is defined as the range of the unique conditional expectations $\tilde{E}_{s}:\tilde{\mathcal{M}}\to\tilde{\mathcal{M}}$ compatible with these weights:
@@
This analytic structure guarantees that the path connects the boundaries smoothly and that the interpolated weights are uniquely determined by the non-commutative $L^{p}$-geometry (identifying $s=1/p$).
Heuristically, this construction signifies that $\tilde{\phi}_{s}$ acts as a non-commutative geometric mean between the reference weight $\tilde{\phi}_{0}$ and the coarse-grained weight $\tilde{\phi}_{1}$. Although formal products of weights are not defined in the operator algebra, we may conceptually visualize this interpolation as:
@@
This geometric averaging property ensures that the path $\tilde{\mathcal{N}}(s)$ follows the geodesic of the underlying modular structure, strictly minimizing the information distance (relative entropy) between the full algebra and the subalgebra at each step.
We now formulate the fundamental physical properties of this path, analogous to the requirements examined for Path A.
Theorem 3.5 (Properties of the Canonical Interpolation Path).
@@
    Nesting (Monotonicity): For any $0\leq s\leq s^{\prime}\leq 1$, the inclusion relation $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ holds. This reflects the monotonic coarse-graining of information, structurally analogous to a renormalization group flow.
@@
These properties are direct consequences of the analytic structure of the Haagerup-Kosaki construction. Unlike naive spectral truncations, this path is generated by the spectral resolution of the Radon-Nikodym derivative $h$, which ensures structural continuity via the modular automorphism group.
(1) Boundary Conditions: The path relies on the analytic continuation of the perturbed weights. At $s=0$, the generator $h^{0}=\mathbf{1}$ acts trivially, implying that the associated projection is the identity map ($\tilde{E}_{0}=\text{id}_{\tilde{\mathcal{M}}}$). Conversely, at $s=1$, the construction converges to the operator $h^{1}$ which generates the target conditional expectation $\tilde{E}$. Thus, $\tilde{E}_{1}$ recovers the original projection onto $\tilde{\mathcal{N}}$, eliminating the discontinuity at $s=0$ characteristic of spectral cut methods.
(2) Nesting: This property follows from the spectral ordering of the Radon-Nikodym derivative $h$. The parameter $s$ effectively controls the resolution of the algebra relative to the reference state. Since $h$ is a positive operator, the family of spectral projections associated with $h^{s}$ defines a monotonic hierarchy of subspaces. Mathematically, increasing $s$ corresponds to a stronger restriction on the algebra, strictly implying $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ for $s^{\prime}\geq s$.
(3) Idempotency: This follows immediately from the nesting property $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$. By the tower property of conditional expectations, the composition of mappings onto nested subalgebras reduces to the mapping onto the smaller subalgebra, yielding $\tilde{E}_{s^{\prime}}\circ\tilde{E}_{s}=\tilde{E}_{s^{\prime}}$. ∎
The canonical $L^{p}$-interpolation provides more than mathematical smoothness; it imposes a strict ordering on information. The nesting property $\tilde{\mathcal{N}}(s^{\prime})\subseteq\tilde{\mathcal{N}}(s)$ established in Theorem 3.5 identifies $s$ as a coarse-graining scale: increasing $s$ corresponds to systematically discarding information, strictly analogous to the irreversible flow of a renormalization group. This structure now enables us to construct the unitary operator $\tilde{U}(s):=J_{\tilde{\mathcal{M}}}J_{\tilde{\mathcal{N}}(s)}$, which serves as the lifted counterpart to the physical recovery path $U(s)$ originally envisioned in Section 3.1. While the original path $U(s)$ was ill-defined due to the Type III obstruction, this lifted operator $\tilde{U}(s)$ is rigorously defined within the crossed-product framework. It acts as an information recovery process, reversing the coarse-graining by translating information from the finer-grained algebra back into the decodable sector. Thus, this construction not only connects the two algebras but also establishes the rigorous framework required to define the unitary recovery path, which we explicitly construct in the following section.